OpenStudy (anonymous):

A sphere has a radius of 4x + 4. Which polynomial in standard form best describes the total surface area of the sphere? Use the formula S = 4r2 for the surface area of a sphere.

OpenStudy (anonymous):

The surface area is \(S = 4\pi r^2\) as you are given and you know that the radius is \(r = 4x+4\). If you substitute that into the equation and simplify what do you get? \[S = 4\pi(4x+4)^2=\cdots \]

OpenStudy (anonymous):

A. 64pix^2 + 128pix + 64pi B. 64pix^2 – 128pix + 64pi C. 64pix^2 + 128pix – 64pi D. 64pix^2 – 128pix – 64pi

OpenStudy (anonymous):

I am not going to tell you the answer. Please attempt it for yourself.

OpenStudy (zehanz):

You could save a little on the calculations by factoring out 4: S=4π(4x+4)²=4π((4(x+1))²=4π16(x+1)²=64π(x+1)²=...

OpenStudy (zehanz):

@yakeyglee: don't be too harsh ;)

OpenStudy (anonymous):

I just need help ! I don't understand how to do it

OpenStudy (anonymous):

I'm new to this, I just created me an account like really

OpenStudy (anonymous):

What is not clear about our explanations? They are guiding you to the answer.

OpenStudy (zehanz):

AFAIK, you now have 64π(x+1)², so first expand (x+1)²=(x+1)(x+1)=... YOu must have learned how to do that (FOIL?)

OpenStudy (anonymous):

Can you just please give me the answer?

OpenStudy (annipuppi):

I have this same question on my homework... I'm really confused on it. so far after reading this I have 64pi(x^2+2x+2) am I doing it right so far? What do I do next? @yakeyglee @ZeHanz

OpenStudy (zehanz):

You're doing fine! Now you just have to expand the brackets (use the distributive property), so multiply 64pi with x², with 2x and with 2...

OpenStudy (annipuppi):

so 64pix^2*128x*128?

OpenStudy (zehanz):

That would be 64pi(x^2+2x+1) = 64pix^2+128pix+64pi